EXERCISE 10–2 Direct Labor Variances LO10–2
SkyChefs, Inc., prepares in-flight meals for a number of major airlines. One of the company’s products is grilled salmon in dill sauce with baby new potatoes and spring vegetables. During the most recent week, the company prepared 4,000 of these meals using 960 direct labor-hours. The company paid its direct labor workers a total of $19,200 for this work, or $20.00 per hour.

According to the standard cost card for this meal, it should require 0.25 direct labor-hours at a cost of $19.75 per hour.

Required:

1. What is the standard labor-hours allowed (SH) to prepare 4,000 meals?
2. What is the standard labor cost allowed (SH × SR) to prepare 4,000 meals? 3. What is the labor spending variance?
4. What is the labor rate variance and the labor efficiency variance?
Answer #1
Ques 1
standard labor hours allowed = 4000 meals * 0.25 Direct hours =          1,000 hours
Ques 2
standard labor cost = 1000 hours * $19.75= $        19,750
Ques 3
Spending varaince = AH* AR - SH*SR
Spending varaince = 19200-19750= $              550 Favourable
Ques 4
Labor rate variance= AH* (AR-SR)=960*(20-19.75)
Labor rate variance= AH* (AR-SR)= 240 Unfavourable
Labor efficiency variance = SR* (AH-SH)
Labor efficiency variance = 19.75* (960-1000)
Labor efficiency variance = 790 Favourable